OpenStudy (anonymous):

Which fraction will simplify to 7/8 ? A. 1/3 B. 8/9 C. 21/24 D. 28/40

OpenStudy (anonymous):

3*8=?

OpenStudy (emilylauren):

C

OpenStudy (anonymous):

24

OpenStudy (emilylauren):

21/3=7 24/3=8 Simplified: 7/8

OpenStudy (anonymous):

A fraction is what is known as a ratio a:b. Basically we divide the number a by the number b and deem it a fraction. Fractions have many properties to them. One common property is that fractions can be simplified by dividing out common factors. For example: 25/20 and 5/4 Notice how 5*5=25 and 4*5=20, both have a common factor of 5. All I am doing is dividing a 5 out from both parts. 5(5)/4(5)=25/20. Since 5/5=1, I am left with just 5/4=25/20. Now try the problem up above and tell me what you think.